# Revealing the Origin of Neutrino Masses through Displaced Shower Searches in the CMS Muon System

### v1

@@ V Conclusion
The origin of neutrino masses remains an open issue in particle physics. Although the seesaw mechanism provides an elegant solution, the right-handed neutrinos (RHNs) it predicts are difficult to probe at colliders. This is because the active-sterile mixing is required to be small to generate light neutrino masses in the canonical seesaw. We instead consider that Majorana RHN masses are generated by the spontaneous symmetry breaking of a $B-L$ gauge symmetry, under which the RHNs are charged. The RHNs can then be produced abundantly at colliders but remain long-lived due to the small active-sterile mixing strengths.
A search for the LLPs has been performed at CMS, using the displaced shower signature in its muon system Tumasyan et al. (2021). We have reinterpreted it as a search for RHNs in the $B-L$ gauge model, produced via $pp\to Z^{\prime},Z,h,\Phi\to NN$ and decaying at displaced vertices in the CMS endcap. We found that by applying the same trigger and reconstruction, the Higgs portals $h$ and $\Phi$ can be used to probe RHNs with masses $m_{N}\approx 40$ GeV and active-sterile mixing strengths $|V_{\ell N}|^{2}\approx 10^{-12}$ using existing CMS data and for $B-L$ model parameters satisfying current constraints. The gauge portals $Z$ and $Z^{\prime}$ offer less promise in the near future.
Our results illustrate that searches for displaced showers in the CMS muon system are powerful probes to reveal the origin of neutrino masses. Existing CMS data is already excluding yet unexplored parameter space in the well-motivated $B-L$ gauge model that covers the canonical seesaw floor, i.e., where a light neutrino mass scale of the order $10^{-2}~{}\text{eV}\lesssim m_{\nu}\lesssim 0.5$ eV is generated. The work motivates a dedicated search for displaced showers in the muon system with potential optimizations applied, such as for 3-body decay modes relevant for our RHN signatures.
@@ Appendix A The Minimal $B-L$ Gauge Model
We here briefly summarize the minimal $B-L$ gauge model, and the constraints on relevant model parameters. The $B-L$ model extends the SM gauge group with the additional $B-L$ symmetry, $SU(3)_{c}\times SU(2)_{L}\times U(1)_{Y}\times U(1)_{B-L}$. The particle content is also extended, by including the $B-L$ gauge boson $Z^{\prime}$, three fermionic Weyl RHNs $\nu_{R}^{i}$ and the $B-L$ Higgs $\Phi$. The relevant Lagrangian in the interaction eigenstates is
@@
$\phantom{=}-\frac{\epsilon}{2}B^{\mu\nu}Z^{\prime}_{\mu\nu}\left\{+m^{2}_{ZZ^{\prime}}B^{\mu}Z^{\prime}_{\mu}\right\}$
[…]
$\phantom{=}-\sum_{\alpha,j}\left(\lambda_{D}^{\alpha j}\bar{L}^{\alpha}\cdot\tilde{H}\nu_{R}^{j}+\text{h.c.}\right)-{\cal V}(H,\Phi),$ (A.1)
with the field strength tensor of the $B-L$ gauge group $Z^{\prime}_{\mu\nu}=\partial_{\mu}Z^{\prime}_{\nu}-\partial_{\nu}Z^{\prime}_{\mu}$, the hypercharge, $B_{\mu\nu}=\partial_{\mu}B_{\nu}-\partial_{\nu}B_{\mu}$, and the covariant derivative $D_{\mu}=D_{\mu}^{\text{SM}}-ig_{B-L}Q_{B-L}Z_{\mu}^{\prime}$, including the $B-L$ contribution, with $g_{B-L}$ and $Q_{B-L}$ being the $B-L$ gauge coupling and charge, respectively. In Eq. (A), $H$ is the SM Higgs doublet, with $\tilde{H}=i\sigma^{2}H^{*}$ and $L^{\alpha}$ are the SM lepton doublets. The SM particles have their canonical $Q_{B-L}$ charges while that of the exotic particles take the values $Q_{B-L}(Z^{\prime})=0$, $Q_{B-L}(\nu_{Ri})=-1$ and $Q_{B-L}(\Phi)=2$, while they are singlets under the SM gauge symmetries.
@@ A.1 Higgs Sector
${\cal V}(H,\Phi)$ $=m^{2}H^{\dagger}H+\mu^{2}|\Phi|^{2}+\lambda_{1}(H^{\dagger}H)^{2}$
$+\lambda_{2}|\Phi|^{4}+\lambda_{3}H^{\dagger}H|\Phi|^{2}.$ (A.2)
Both the SM and $B-L$ Higgs acquire vacuum expectation values, $v=\langle H^{0}\rangle$, $v_{B-L}=\langle\Phi\rangle$, leading to a breaking of the model’s gauge group to $SU(3)_{c}\times U(1)_{\text{EM}}$. In addition, the scalar potential in Eq. (A.1) gives rise to the mass matrix of the Higgs fields $(H^{0},\Phi)$ at tree level Robens and Stefaniak (2015),
@@
$M^{2}_{h,\Phi}$ $=\lambda_{1}v^{2}+\lambda_{2}v_{B-L}^{2}$
$-(+)\sqrt{(\lambda_{1}v^{2}-\lambda_{2}v_{B-L}^{2})^{2}+(\lambda_{3}v_{B-L}v)^{2}},$ (A.4)
@@
where the mixing angle $\alpha$ is given through
@@
We take $m_{h}=m_{H^{0}}\approx 125~{}\text{GeV}<m_{\Phi}$ where we drop the prime on the mass eigenstate [^1] The case where $m_{\Phi}<m_{h}$ is considered in Liu et al. (2022c)..
Figure 3: Current limits on the Higgs mixing strength $\sin\alpha$, adapted from Robens (2023). The bound from measuring the Higgs signal rate has been updated using the LHC Run 2 results Robens (2023); Papaefstathiou et al. (2022); ATLAS (2021). Their projection at the HL-LHC with 3000 fb-1 integrated luminosity, taking the uncertainty $\delta\mu=|1-\sigma(pp\to h)/\sigma(pp\to h)_{\text{SM}}|=0.02$, is shown for comparison Cepeda et al. (2019); ATLAS and CMS (2019).
A summary of recent experimental limits on the Higgs mixing within an exotic singlet scalar can be found in Robens (2023). The Higgs mixing can be probed by direct searches for a heavy scalar and a measurement of the Higgs signal rate Robens and Stefaniak (2015); Robens (2021, 2023). The existence of an additional scalar also introduces a shift to the $W$ boson mass López-Val and Robens (2014); Papaefstathiou et al. (2022). A limit from the $W$ boson mass is derived by comparing the experimental Particle Data Group value, $m_{W}^{\text{exp}}=80.379\pm 0.012$ GeV and the SM predicted value, $m_{W}^{\text{SM}}=80.356$ GeV Workman et al. (2022). Lastly, requiring the scalar coupling $\lambda_{1}$ to remain perturbative can also be used to infer a limit on the mixing. From Ref. Robens (2023); Papaefstathiou et al. (2022) using LHC Run 2 results ATLAS (2021), an upper limit from the signal rate can be obtained at $\sin\alpha\lesssim 0.24$. At the 14 TeV HL-LHC with 3000 fb-1 integrated luminosity, a measurement of the Higgs signal rate is expected to achieve an uncertainty of $\delta\mu=|1-\sigma(pp\to h)/\sigma(pp\to h)_{\text{SM}}|=0.02$ Cepeda et al. (2019); ATLAS and CMS (2019), with a projected sensitivity on the Higgs mixing of $\sin\alpha\approx 0.14$.
@@
Besides SM Higgs production at the LHC, we also consider the production of the $B-L$ Higgs $\Phi$ with subsequent decay to RHNs. The branching ratios of $\Phi$ to relevant decay products is shown in Fig. 4 as a function of its mass, calculated in MadGraph for the $B-L$ gauge model. The $B-L$ Higgs decays to SM final states as well as a pair of RHNs $\Phi\to NN$ where we take the RHN mass $m_{N}=0.3\times m_{\Phi}$ that approximately maximizes the partial decay rate. The decays to SM particles are suppressed by the Higgs mixing which is taken at a small value $\sin\alpha=0.08$ in the figure to illustrate the potential dominance of the decay $\Phi\to NN$ for $m_{\Phi}\lesssim 160$ GeV. The $B-L$ Higgs VEV is fixed at $\langle\Phi\rangle=3.75$ TeV. As noted in the main section, as long as $\Phi\to NN$ is not dominant, the cross section $\sigma(pp\to\Phi\to NN)$ is largely independent of the Higgs mixing as it cancels out between production and decay. As can be seen, Br$(\Phi\to NN)\approx 0.5-0.85$ is large for $m_{\Phi}\lesssim 2m_{W}\approx 160$ GeV.
@@ A.2 Gauge Sector
Figure 5: Current limits on the $B-L$ gauge coupling $g_{B-L}$ (left) and the $B-L$ VEV $\langle\Phi\rangle=m_{Z^{\prime}}/(2g_{B-L})$ (right) as a function of the $Z^{\prime}$ mass $m_{Z^{\prime}}$ from resonance searches at CMS CMS (2019); Sirunyan et al. (2020, 2021); Chiang et al. (2019); Liu et al. (2022b), ATLAS Aad et al. (2019b); Chiang et al. (2019), LHCb Aaij et al. (2020), electroweak precision tests Cacciapaglia et al. (2006); Alcaraz et al. (2006) and searches for invisible $Z^{\prime}$ final states Fox et al. (2011) (LEP). Also shown are the projected sensitivities at ATLAS and CMS for the 3000 fb-1 HL-LHC.
@@
Arguably the most direct way to probe the $B-L$ model is through extra gauge boson with mass $m_{Z^{\prime}}=2g_{B-L}v_{B-L}$. In Fig. 5 (left), the current limits and projected sensitivities on the gauge coupling $g_{B-L}$ as a function of $m_{Z^{\prime}}$ are shown. For the parameter space of our interests, $1010~{}\text{GeV}<m_{Z^{\prime}}<$ 10 TeV, the existing limits mainly arise from colliders through searches at CMS, ATLAS and LHCb, as well as EW precision tests and searches of invisible final states of $Z^{\prime}$. Limits from the LHCb were originally presented for a dark photon but can be reinterpreted in the $B-L$ model, following Ilten et al. (2018). The limits from EW precision tests (LEP) are effectively on the vacuum expectation value of the $B-L$ Higgs with $v_{B-L}=m_{Z^{\prime}}/(2g_{B-L})\gtrsim 3.5$ TeV  Cacciapaglia et al. (2006); Alcaraz et al. (2006). For light $Z^{\prime}$ with $10~{}\text{GeV}<m_{Z^{\prime}}<m_{W}$, limits are from resonance searches for di-leptons final states at LHCb Aaij et al. (2020) and CMS CMS (2019); Sirunyan et al. (2020). The current limits are $g_{B-L}\lesssim 6\times 10^{-4}$. When $m_{Z^{\prime}}\sim m_{W,Z}$, there are large backgrounds from the SM gauge bosons, hence the limit from LEP become dominant, except for $m_{Z^{\prime}}\lesssim m_{Z}$ when the search of the invisible decay of $Z^{\prime}$ at LEP Fox et al. (2011) is more stringent. The CMS search CMS (2019); Sirunyan et al. (2020) is effective for $200110~{}\text{GeV}<m_{Z^{\prime}}<$ 200 GeV, with an upper limit of $g_{B-L}\lesssim 10^{-3}$. For heavier $Z^{\prime}$, the limits from high mass resonance searches at CMS Sirunyan et al. (2021) and ATLAS Aad et al. (2019b), as well as LEP apply. Among them, the limits from ATLAS are the most stringent for $m_{Z^{\prime}}\lesssim 6$ TeV. Due to the large mass, the background from the SM gauge bosons is suppressed and the limits become less stringent as $Z^{\prime}$ becomes heavier, reaching $g_{B-L}\lesssim 0.2$ when $m_{Z^{\prime}}\approx 6$ TeV. For even heavier $Z^{\prime}$, one can only rely on the limits from LEP. We also show the estimated sensitivities of ATLAS and CMS at the HL-LHC with 3000 fb-1, by scaling the current bounds with luminosity as $g_{B-L}\propto\text{luminosity}^{-1/4}$. The same constraints are displayed in Fig. 5 (right), but interpreted with respect to the $B-L$ VEV $v_{B-L}=\langle\Phi\rangle=m_{Z^{\prime}}/(2g_{B-L})$.
@@
In the minimal $B-L$ model, no kinetic or mass mixing between the SM hypercharge and $B-L$ is considered. Kinetic mixing is incorporated by including the $\epsilon$ term in Eq. (A) Dev et al. (2021). The mass mixing term $m_{ZZ^{\prime}}^{2}$ violates the model’s gauge symmetry but may still be present if the $B-L$ and electroweak symmetry breaking are connected in a broader framework. For our purposes, we take $\epsilon$ and $m_{ZZ^{\prime}}^{2}$ as effective, independent parameters at the electroweak scale. The mass mixing induces a field mixing angle $\gamma$ between the two gauge bosons,
@@
and both kinetic and mass mixing lead to a coupling of the SM-like gauge boson to the $B-L$ current $J_{B-L}^{\mu}$ Lindner et al. (2018), $-\sin\theta_{ZZ^{\prime}}Z_{\mu}J_{B-L}^{\mu}$, with the overall mixing angle

### v2

@@ V Conclusion
The origin of neutrino masses remains an open issue in particle physics. Although the seesaw mechanism provides an elegant solution, the RHNs it predicts are difficult to probe at colliders due to the small active-sterile mixing required to generate light neutrino masses. We instead consider that RHNs are produced via Higgs or gauge portals in New Physics scenarios, with the $B-L$ gauge model as a prototype. As the SM Higgs portal is the most promising channel, we also interpret our results in the $\nu_{R}$. In such scenarios, RHNs can be produced abundantly at colliders but remain long-lived due to the small active-sterile mixing strengths.
We have reinterpreted the CMS analysis [75] as a search for RHNs, produced via $pp\to Z^{\prime},Z,h,\Phi\to NN$ and decaying at displaced vertices in the CMS endcap. We found that the Higgs portals $h$ and $\Phi$ can be used to probe RHNs with masses $m_{N}\approx 40$ GeV and active-sterile mixing strengths $|V_{\ell N}|^{2}\approx 10^{-12}$ using existing CMS data, and for $B-L$ model and $\nu_{R}$ parameters satisfying current constraints. The gauge portals $Z$ and $Z^{\prime}$ offer less promise in the near future.
Our results illustrate that searches for displaced showers in the CMS muon system are powerful probes to reveal the origin of neutrino masses. Existing CMS data is already excluding yet unexplored parameter space in the well-motivated $B-L$ model that covers the seesaw floor, i.e., where a light neutrino mass scale of the order $9\times 10^{-3}~{}\text{eV}\lesssim m_{\nu}\lesssim 0.45$ eV is generated. This can be further generalized in the broad context of the $\nu_{R}$, where New Physics scales $\Lambda_{\text{NP}}\approx 200$ TeV are being tested, which is about one order of magnitude better than the current limit from the Higgs signal strength [108, 109], and is expected to improve to $\Lambda_{\text{NP}}\approx 3000$ TeV at the HL-LHC. Our work motivates a dedicated search for displaced showers in the muon system with potential optimizations applied, such as for 3-body decay modes relevant for our RHN signatures.
@@ Appendix A The Minimal $B-L$ Gauge Model
We here briefly summarize the minimal $B-L$ gauge model, and the constraints on relevant model parameters. The $B-L$ model extends the SM gauge group with an additional Abelian gauge symmetry associated with the $B-L$ quantum number, $SU(3)_{c}\times SU(2)_{L}\times U(1)_{Y}\times U(1)_{B-L}$. The particle content is also extended, by including the $B-L$ gauge boson $Z^{\prime}$, three Weyl RHNs $\nu_{R}^{i}$ and the $B-L$ Higgs $\Phi$. The relevant Lagrangian in the interaction eigenstates is
@@
$\quad-\frac{1}{2}\sum_{i,j}\left(\lambda_{N}^{ij}\bar{\nu}_{R}^{i,c}\Phi\nu_{R}^{j}+\text{h.c.}\right)-\sum_{\alpha,j}\left(\lambda_{D}^{\alpha j}\bar{L}^{\alpha}\cdot\tilde{H}\nu_{R}^{j}+\text{h.c.}\right)-{\cal V}(H,\Phi),$ (A.1)
with the field strength tensor of the $B-L$ gauge group, $Z^{\prime}_{\mu\nu}=\partial_{\mu}Z^{\prime}_{\nu}-\partial_{\nu}Z^{\prime}_{\mu}$, hypercharge, $B_{\mu\nu}=\partial_{\mu}B_{\nu}-\partial_{\nu}B_{\mu}$, and the covariant derivative $D_{\mu}=D_{\mu}^{\text{SM}}-ig_{B-L}Q_{B-L}Z_{\mu}^{\prime}$, including the $B-L$ contribution, with $g_{B-L}$ and $Q_{B-L}$ being the $B-L$ gauge coupling and charge, respectively. In Eq. (A), $H$ is the SM Higgs doublet, with $\tilde{H}=i\sigma^{2}H^{*}$ and $L^{\alpha}$ are the SM lepton doublets. The SM particles have their canonical $Q_{B-L}$ charges while that of the exotic particles take the values $Q_{B-L}(Z^{\prime})=0$, $Q_{B-L}(\nu_{Ri})=-1$ and $Q_{B-L}(\Phi)=2$. The exotic particles are singlets under the SM gauge symmetries.
@@ A.1 Higgs Sector
${\cal V}(H,\Phi)=m^{2}(H^{\dagger}\cdot H)+\mu^{2}|\Phi|^{2}+\lambda_{1}(H^{\dagger}\cdot H)^{2}+\lambda_{2}|\Phi|^{4}+\lambda_{3}(H^{\dagger}\cdot H)|\Phi|^{2}.$ (A.2)
Both the SM and $B-L$ Higgs acquire vacuum expectation values, $v=\langle H^{0}\rangle$, $v_{B-L}=\langle\Phi\rangle$, leading to a breaking of the model’s gauge group to $SU(3)_{c}\times U(1)_{\text{EM}}$. In addition, the scalar potential in Eq. (A.2) gives rise to the mass matrix of the Higgs fields $(H^{0},\Phi)$ at tree level [110],
@@
$M^{2}_{h(\Phi)}$ $=\lambda_{1}v^{2}+\lambda_{2}v_{B-L}^{2}-(+)\sqrt{(\lambda_{1}v^{2}-\lambda_{2}v_{B-L}^{2})^{2}+(\lambda_{3}v_{B-L}v)^{2}},$ (A.4)
@@
where the mixing angle $\alpha$ is determined by
@@
We take $m_{h}\approx 125~{}\text{GeV}<m_{\Phi}$ where we drop the prime on the mass eigenstate  [^1] The case where $m_{\Phi}<m_{h}$ is considered in [70]..
Figure 4: Current limits on the Higgs mixing strength $\sin\alpha$, adapted from [112]. The bound from measuring the Higgs signal rate has been updated using the LHC Run-2 results [112, 113, 114]. Their projection at the HL-LHC with 3000 fb-1 integrated luminosity, taking the uncertainty $\delta\mu=|1-\sigma(pp\to h)/\sigma(pp\to h)_{\text{SM}}|=0.02$, is shown for comparison [105, 106].
A summary of recent experimental limits on the Higgs mixing with an exotic singlet scalar can be found in [112]. The Higgs mixing can be probed by direct searches for a heavy scalar and a measurement of the Higgs signal rate [110, 115, 112]. The existence of an additional scalar also introduces a shift to the $W$ boson mass [116, 113], from which a limit is derived by comparing the experimental Particle Data Group value, $m_{W}^{\text{exp}}=80.379\pm 0.012$ GeV and the SM predicted value, $m_{W}^{\text{SM}}=80.356$ GeV [117]. Lastly, requiring the scalar coupling $\lambda_{1}$ to remain perturbative can also be used to infer a limit on the mixing. From Refs. [112, 113], using LHC Run-2 results [114], an upper limit from the signal rate can be obtained, $\sin\alpha\lesssim 0.24$. At the 14 TeV HL-LHC with 3000 fb-1 integrated luminosity, a measurement of the Higgs signal rate is expected to achieve an uncertainty of $\delta\mu=|1-\sigma(pp\to h)/\sigma(pp\to h)_{\text{SM}}|=0.02$ [105, 106], with a projected sensitivity on the Higgs mixing of $\sin\alpha\approx 0.14$. The current constraints and the future HL-LHC sensitivity are shown in Fig. 4.
@@
Besides SM Higgs production at the LHC, we also consider the production of the $B-L$ Higgs $\Phi$ with subsequent decay to RHNs. The branching ratios of $\Phi$ to relevant decay products are shown in Fig. 5 as a function of its mass, calculated in MadGraph for the $B-L$ gauge model. The $B-L$ Higgs decays to SM final states as well as a pair of RHNs, $\Phi\to NN$, where we take the RHN mass $m_{N}=0.3\times m_{\Phi}$ which approximately maximizes the partial decay rate. The decays to SM particles are suppressed by the Higgs mixing which is taken at a small value $\sin\alpha=0.08$ in the figure to illustrate the potential dominance of the decay $\Phi\to NN$ for $m_{\Phi}\lesssim 160$ GeV. The $B-L$ Higgs VEV is fixed at $\langle\Phi\rangle=3.75$ TeV. As noted in the main text, as long as $\Phi\to NN$ is not dominant, the cross section $\sigma(pp\to\Phi\to NN)$ is largely independent of the Higgs mixing as it cancels out between production and decay. As can be seen, Br$(\Phi\to NN)\approx 0.5-0.85$ is large for $m_{\Phi}\lesssim 2m_{W}\approx 160$ GeV.
@@ A.2 Gauge Sector
Figure 6: Current limits on the $B-L$ gauge coupling $g_{B-L}$ (left) and the $B-L$ VEV $\langle\Phi\rangle=m_{Z^{\prime}}/(2g_{B-L})$ (right) as a function of the $Z^{\prime}$ mass $m_{Z^{\prime}}$ from resonance searches at CMS [118, 119, 57, 60], ATLAS [80, 57], LHCb [120], electroweak precision tests [86, 87] and searches for invisible $Z^{\prime}$ final states [121] (LEP). Also shown are the projected sensitivities at ATLAS and CMS for the 3000 fb-1 HL-LHC.
@@
Arguably the most direct way to probe the $B-L$ model is through the extra gauge boson with mass $m_{Z^{\prime}}=2g_{B-L}v_{B-L}$. In Fig. 6 (left), the current limits and projected sensitivities on the gauge coupling $g_{B-L}$ as a function of $m_{Z^{\prime}}$ are shown. For the parameter space of our interest, $1010~{}\text{GeV}<m_{Z^{\prime}}<$ 10 TeV, the existing limits mainly arise from colliders through searches at CMS, ATLAS and LHCb, as well as EW precision tests and searches of invisible final states of the $Z^{\prime}$. Limits from the LHCb were originally presented for a dark photon but can be reinterpreted in the $B-L$ model, following [122]. The limits from EW precision tests (LEP) are effectively on the vacuum expectation value of the $B-L$ Higgs with $v_{B-L}=m_{Z^{\prime}}/(2g_{B-L})\gtrsim 3.5$ TeV  [86, 87]. For light $Z^{\prime}$ with $10~{}\text{GeV}<m_{Z^{\prime}}<m_{W}$, limits are from resonance searches for di-leptons final states at LHCb [120] and CMS [118]. The current best limits are $g_{B-L}\lesssim 6\times 10^{-4}$ for $m_{Z^{\prime}}\lesssim 70$ GeV. When $m_{Z^{\prime}}\sim m_{W,Z}$, there are large backgrounds from the SM gauge bosons, hence the limit from LEP becomes dominant, except for $m_{Z^{\prime}}\lesssim m_{Z}$ when the search of the invisible decay of $Z^{\prime}$ at LEP [121] is more stringent. The CMS search [118] is effective for $200110~{}\text{GeV}<m_{Z^{\prime}}<$ 200 GeV, with an upper limit of $g_{B-L}\lesssim 10^{-3}$. For heavier $Z^{\prime}$, the limits from high mass resonance searches at CMS [119] and ATLAS [80], as well as LEP apply. Among them, the limits from ATLAS are the most stringent for $m_{Z^{\prime}}\lesssim 6$ TeV. Due to the large mass, the background from the SM gauge bosons is suppressed and the limits become less stringent as $Z^{\prime}$ becomes heavier, reaching $g_{B-L}\lesssim 0.2$ when $m_{Z^{\prime}}\approx 6$ TeV. For even heavier $Z^{\prime}$, one can only rely on the limits from LEP. We also show the estimated sensitivities of ATLAS and CMS at the HL-LHC with 3000 fb-1, by scaling the current bounds with luminosity as $g_{B-L}\propto\text{luminosity}^{-1/4}$. The same constraints are displayed in Fig. 6 (right), but interpreted with respect to the $B-L$ VEV $v_{B-L}=\langle\Phi\rangle=m_{Z^{\prime}}/(2g_{B-L})$.
@@
In the minimal $B-L$ model, no kinetic or mass mixing between the SM hypercharge and $B-L$ is considered. Kinetic mixing is incorporated by including the $\epsilon$ term in Eq. (A) [82]. The mass mixing term $m_{ZZ^{\prime}}^{2}$ violates the model’s gauge symmetry but may still be present if the $B-L$ and electroweak symmetry breaking are connected in a broader framework. For our purposes, we take $\epsilon$ and $m_{ZZ^{\prime}}^{2}$ as effective, independent parameters at the electroweak scale. The mass mixing induces a field mixing angle $\gamma$ between the two gauge bosons,
@@
and both kinetic and mass mixing lead to a coupling of the SM-like gauge boson to the $B-L$ current $J_{B-L}^{\mu}$ [123], $-\sin\theta_{ZZ^{\prime}}Z_{\mu}J_{B-L}^{\mu}$, with the overall mixing angle